# CERN-TH-2024-065 ET-0343A-24

### v1

@@ II.1 SGWB detection methods
namely the sum of the squared SNR of each detector pair Hazboun et al. (2019). Note that the assumption of uncorrelated noise in $N$ detectors applies to PTA experiments [^3] This assumption applies well to intrinsic pulsar noises. However, some noise sources, such as clock, planetary ephemeris errors or poorly modeled solar wind, could imprint common correlated noises in pulsar data Tiburzi et al. (2015)., where each pulsar can be considered as a detector. It does not apply to the three co-located interferometers composing the Einstein Telescope, which we treat separately below. Note also that the SNR scales with the square root of the observing time $T_{\mathrm{obs}}$.
@@ II.2.1 Earth-based detectors
Regarding ET, we work with the current baseline configuration consisting of a triangular arrangement of three Michelson interferometers with 10km arms, situated underground to minimize seismic noise. Since the three V-shape interferometers are co-located, one cannot actually use Eq. (11), as the noises will be partially correlated among the three interferometers. To take into account these partial correlations, we follow a similar approach to Branchesi et al. (2023) to compute the effective strain noise PSD. Note, however, that our resulting sensitivity differs slightly from the one calculated in Branchesi et al. (2023), see Appendix B for more details. In brief, we take the noise power spectral density $S_{\mathrm{ET-L}}$ for one ET L-shape interferometer from ET- , including a factor of $\sin^{2}(\pi/3)$ due to the triangular shape. We then use an excess cross-correlation power statistic between the three interferometers pairs, assuming a 20% correlation amongst interferometers as in Branchesi et al. (2023). In the weak signal regime (see Appendix B), keeping only the linear terms in the noise correlation coefficient $\Gamma^{(n)}_{12}=0.2$ (see Eq. 131), we then find
@@ II.2.2 Space-based detector
The data employed for GW search in LISA will be obtained from Time Delay Interferometry (TDI). This mitigates laser frequency noise by delaying and recombining individual measurements along spacecraft links (for further details, see Prince et al. (2002); Tinto and Dhurandhar (2021)). This process yields three Michelson TDI channels, often denoted as $X$, $Y$, and $Z$, obtained by combining measurements in each pair of arms. Assuming that the length of the arms is constant in time, it is then possible to construct uncorrelated combinations of these variables which are the $A$, $E$ and $T$ channels Estabrook et al. (2000); Vallisneri and Galley (2012). Clearly one cannot cross-correlate the output from these three channels to search for an isotropic SGWB, as the signal as well as the noise are uncorrelated amongst the channels. Furthermore, the $T$ combination, which corresponds to a Sagnac interferometer, is insensitive to the GW signal at low frequency Hogan and Bender (2001). It has been proposed to use the auto-correlation of the $A$ and $E$ channels to look for excess power due to the presence of a SGWB. [^5] Many other methods are proposed in the literature, see Boileau et al. (2021) and references therein. As explained earlier, this method is not optimal since it assumes that we have a perfect estimation of the noise in the $A$ and $E$ channels. Thanks to the uncorrelated nature of these two combinations, one obtains for the total GWB SNR in LISA
@@
The detection of cosmological GWB with LISA will be affected by astrophysical foregrounds. The first we consider originates from the galactic population of white dwarf binaries. An estimation of the resulting foreground PSD after (perfect) substraction of individual bright binaries can be found in Babak et al. (2021); Karnesis et al. (2021). The level of this noise is expected to exceed the detector noise around 1mHz. Consequently, we combine the PSD of this foreground with that of the detector in quadrature. The second foreground we consider is due to the extra galactic population of unresolvable white dwarf binaries and stellar-mass black hole (BH) binaries. We model the former using Eq. (25) of Staelens and Nelemans (2024) and the latter using Eq. (20), with $\Omega_{A}=7.87\cdot 10^{-13}$ and $f_{*}=0.003$ Hz Babak et al. (2023). In the same way as for the Earth-based detectors, we compute the SNR of this foreground using Eq. (21). We find $\mathrm{SNR}^{\mathrm{LISA}}_{\mathrm{BBH+WDB}}\simeq 314$, which is dominated by the population of extra galactic WDB since $\mathrm{SNR}^{\mathrm{LISA}}_{\mathrm{BBH}}\simeq 53$ in agreement with Pieroni and Barausse (2020).
@@ II.2.3 Pulsar Timing Arrays
Regarding the computation of the effective strain noise PSD of SKA, we employ Eq. (11) using the hasasia Python package to produce realistic sensitivity curves Hazboun et al. (2019). Indeed, PTA experiments are not only limited by the white noise from the radio telescopes and pulse jitter Goncharov et al. (2020), but also by three main components: (i) the uncertainty in the timing model parameters used to construct the expected times of arrival of the radio pulses Van Haasteren et al. (2009); Blandford et al. (1984), (ii) the potential intrinsic red noise that may impact the stability of pulsar spin at low frequencies Shannon and Cordes (2010), and (iii) the chromatic noise induced by electron density fluctuations in the interstellar medium along the path of the radio pulses to Earth Jones et al. (2017); Lam et al. (2016). The latter can, in principle, be mitigated using multi-band observation but it is still anticipated to affect the sensitivity of the experiment Janssen et al. (2015). For (i) we consider the pulsar spin rate, its derivative, its sky position and distance in the timing model, see Hazboun et al. (2019) for their effects on the PTA sensitivity. For (ii) and (iii), using the results from currently observed millisecond pulsars Agazie et al. (2023d); Antoniadis et al. (2023c), we add an intrinsic achromatic red noise component to one-third of the pulsars in the array. We parameterize this noise by a power law PSD with an amplitude (at the reference frequency of 1/year) and spectral index, in terms of characteristic strain, of $\log_{10}A=-13.5$ and $\alpha=1/2$.
@@ III Primordial sources: overview on first order phase transition
During a PT, the order parameter (generally a scalar field in the cosmological case) changes from a metastable vacuum into a more energetically favoured one. If the system evolves maintaining thermal equilibrium, as predicted by the SM for both the QCD Aoki et al. (2006); Stephanov (2006) and the EW PTs Gurtler et al. (1997); Aoki (1997); Kajantie et al. (1996a, b); Laine and Rummukainen (1999), no significant GW emission is expected. While GWs are indeed produced by a plasma in equilibrium, their amplitude is too small to be of observational relevance if the plasma is the one of the SM Ghiglieri et al. (2020), and even when hidden sectors are present Drewes et al. (2023) (note, however, that the question remains open in the context of thermal inflation Klose et al. (2022a, b)). On the other hand, if the PT proceeds via tunnelling or thermal fluctuations over a potential barrier Linde (1983); Hogan (1983); Turner et al. (1992), bubbles of the new phase are nucleated and expand, and ultimately percolate to complete the transition. A SGWB is then sourced by the anisotropic stresses generated by the scalar field configuration when the bubbles collide Kosowsky et al. (1992); Kosowsky and Turner (1993); Caprini et al. (2008); Huber and Konstandin (2008); Jinno and Takimoto (2017); Cutting et al. (2018, 2021), and/or by the bulk fluid motion induced by the bubble expansion and collision Witten (1984); Hogan (1986); Kamionkowski et al. (1994); Hindmarsh et al. (2014, 2015, 2017); Cutting et al. (2020); Hindmarsh and Hijazi (2019); Kosowsky et al. (2002); Dolgov et al. (2002); Caprini and Durrer (2006); Kahniashvili et al. (2008); Caprini et al. (2009b); Brandenburg et al. (2017); Niksa et al. (2018); Roper Pol et al. (2020a).
@@ III.1 Dynamics of a first order phase transition
The starting point to determine the PT dynamics within a given model is the effective potential of the system. In many scenarios the barrier is not present at tree level, but is it induced by radiative and/or finite-temperature corrections. An important source of uncertainty is therefore linked to the fact that a perturbative treatment of the finite temperature potential can be significantly inaccurate. The best resummation techniques are based on matching to a three-dimensional effective field theory, but they can be heavy to implement, and must be carried out on a model-by-model basis (for a review, see Athron et al. (2024)). This uncertainty can severely affect SGWB predictions, via the inaccurate evaluation of the PT parameters entering the GW signal Kainulainen et al. (2019); Gould et al. (2019); Niemi et al. (2019); Croon et al. (2021); Kierkla et al. (2024); Niemi et al. (2021); Schicho et al. (2021); Gould and Tenkanen (2024); Lewicki et al. (2024); Banerjee et al. (2024). Correspondingly, it can also compromise the reconstruction of the model parameters, in case a SGWB signal from a first order PT is detected Bian et al. (2020); Hashino et al. (2017); Friedrich et al. (2022); Caprini et al. (2024).
@@ III.2 Gravitational wave generation from a first order phase transition
It is customary to define the energy fraction $K=\rho_{s}/\rho_{\mathrm{tot}}$, where $\rho_{s}$ denotes the energy density of the GW sourcing process (i.e., kinetic energy in the scalar field and/or in the fluid motion), and $\rho_{\mathrm{tot}}$ the total energy density in the universe at the PT time. The source anisotropic stresses can then be normalised as $\tilde{\Pi}_{ij}=\Pi_{ij}/(K\rho_{\mathrm{tot}})$, with $\Pi_{ij}=T_{ij}^{TT}$ and $T_{ij}$ is the spatial part of the source energy momentum tensor. For the three sourcing processes described above, $\tilde{\Pi}_{ij}$ is a random variable, since the bubble nucleation process is stochastic. $\tilde{\Pi}_{ij}$ also has a finite lifetime $\delta\tau_{\mathrm{fin}}=\tau_{\mathrm{fin}}-\tau_{*}$, where $\tau$ denotes conformal time, we assume radiation domination and set $\tau_{*}=\mathcal{H}_{*}^{-1}=(a_{*}H_{*})^{-1}$ as the initial time of action of the source. $\delta\tau_{\mathrm{fin}}$ can be short compared to the Hubble time (e.g. for bubble collision), but can also be of several Hubble times (e.g. for turbulence).
After the decay of the source, the GWs become freely propagating plane waves, redshifting with the expansion of the universe. The SGWB power spectrum today satisfies then:
@@
$h^{2}\mathrm{\Omega}_{\rm{GW}}(\tau_{0},f)=\frac{h^{2}}{\rho_{c}}\left(\frac{a_{\mathrm{fin}}}{a_{0}}\right)^{4}\rho_{\mathrm{fin}}\,\mathrm{\Omega}_{\rm{GW}}(\tau_{\mathrm{fin}},f)=h^{2}F_{\mathrm{GW},0}\,\mathrm{\Omega}_{\rm{GW}}(\tau_{\mathrm{fin}},f)\,,~{}~{}~{}~{}\mathrm{with}~{}~{}h^{2}F_{\mathrm{GW},0}\simeq 1.64\cdot 10^{-5}\left(\frac{100}{g_{\mathrm{fin}}}\right)^{1/3}.$ (42)
@@
Assuming that the source is operating during the radiation dominated era with $\rho_{\mathrm{tot}}(\eta)\propto a^{-4}(\eta)$, neglecting possible changes in the relativistic degrees of freedom, and averaging over the irrelevant, rapid oscillations in time of the plane waves, the GW energy density fraction when the source has stopped operating becomes (where $k$ denotes the comoving wave-number) Caprini et al. (2008, 2009b); Caprini and Figueroa (2018)
$\mathrm{\Omega}_{\rm{GW}}(\tau_{\mathrm{fin}},k)=\frac{3}{4\pi^{2}}K^{2}k^{3}\int^{\tau_{\mathrm{fin}}}_{\tau_{*}}\frac{\mathrm{d}\eta}{\eta}\int^{\tau_{\mathrm{fin}}}_{\tau_{*}}\frac{\mathrm{d}\zeta}{\zeta}\cos[k(\eta-\zeta)]P_{\tilde{\Pi}}(k,\eta,\zeta)\,,$ (43)
@@
where the GW source $P_{\tilde{\Pi}}(k,\eta,\zeta)$ is the unequal time correlator (UETC) of the anisotropic stress in Fourier space,
$\langle\tilde{\Pi}_{ij}(\mathbf{k},\eta)\,\tilde{\Pi}_{ij}^{*}(\mathbf{q},\zeta)\rangle=(2\pi)^{3}\,\delta^{(3)}(\mathbf{k}-\mathbf{q})\,P_{\tilde{\Pi}}(k,\eta,\zeta)\,.$ (44)
@@
In order to find the SGWB signal from a first order PT, one needs therefore to determine the anisotropic stress UETC of each of the the sourcing processes described above. This is a complicated task, which can be tackled both analytically and numerically. However, it turns out that, by making use of two simple models for the UETC, it is possible to understand the scaling with the sources parameters provided by more refined analytical and numerical evaluations. We present these models below, as an update of what discussed in sections 3.1 and 3.2 of Caprini et al. (2020). In the following we will then clarify the connection between the general scalings provided by these simple models, and what found in specific analytical and numerical evaluations on a source-by-source basis.
@@
  • The first model consists in assuming that the UETC is simply constant in time, $P_{\tilde{\Pi}}(k,\eta,\zeta)=P_{\tilde{\Pi}}(k)$ Roper Pol et al. (2022a). As we will see, the constant-in-time approximation can be applied to bubble collision and to (M)HD turbulence.
@@
  • The second model consists in assuming that the UETC is stationary in time, $P_{\tilde{\Pi}}(k,\eta,\zeta)=f(k,\zeta-\eta)$ Hindmarsh et al. (2014, 2015). This assumption has been used mainly for SGWB generation by sound waves.
    Since the source is necessarily dissipated at some point, it is “stationary” but only within its lifetime $\delta\tau_{\mathrm{fin}}$. Upon changing variable $\tau_{-}=\zeta-\eta$, Eq. 43 takes the form:
    $\mathrm{\Omega}_{\rm{GW}}(\tau_{\mathrm{fin}},k)\simeq\frac{3}{4\pi^{2}}K^{2}k^{3}\left[\int_{\tau_{*}}^{\tau_{\mathrm{fin}}}\frac{\mathrm{d}\eta}{\eta}\int_{\tau_{*}-\eta}^{\tau_{\mathrm{fin}}-\eta}\mathrm{d}\tau_{-}\frac{\cos(k\tau_{-})f(k,\tau_{-})}{\eta+\tau_{-}}\right]\,.$ (49)
@@
    $\mathrm{\Omega}_{\rm{GW}}(\tau_{\mathrm{fin}},k)\simeq\frac{3}{4\pi^{2}}K^{2}k^{3}\int_{\tau_{*}}^{\tau_{\mathrm{fin}}}\frac{\mathrm{d}\eta}{\eta^{2}}\int_{-\infty}^{+\infty}\mathrm{d}\tau_{-}\cos(k\tau_{-})f(k,\tau_{-})\simeq\frac{3}{4\pi^{2}}K^{2}\frac{\mathcal{H}_{*}\delta\tau_{\mathrm{fin}}}{1+\mathcal{H}_{*}\delta\tau_{\mathrm{fin}}}[\mathcal{H}_{*}k^{3}F(k)]\,,$ (50)

### v2

@@ I Introduction on GWs from the early universe
Note: this review has a broad scope, and thus some topics are necessarily treated in less detail than could be done otherwise. Throughout the review, we provide many references, where more details may be found. Here we highlight some books and other reviews which have overlap with the different sections of this review. For general books covering GWs see e.g. Maggiore (2007, 2018); Poisson and Will (2014). Reference e.g. Caprini and Figueroa (2018) is a detailed review on the cosmological SGWB, as well as different sources of that background. SGWB detection methods for different detectors are discussed in e.g. Romano and Cornish (2017); Thrane and Romano (2013); van Remortel et al. (2023). Our review of the physics of first order phase transitions leading to a SGWB can be complemented with e.g. Hindmarsh et al. (2021a); Caldwell et al. (2022); Athron et al. (2024); Roshan and White (2024) for more details. Topological defects are the subject of different books e.g. Vilenkin and Shellard (2000) and reviews e.g. Hindmarsh and Kibble (1995); Vachaspati et al. (2015). The SGWB generated by standard local cosmic strings is discussed in detail in Auclair et al. (2023a). For global cosmic strings and domain walls the situation is more uncertain and there has been renewed interest and progress in recent years (see e.g. Saikawa (2017) for an older review on GWs from domain walls). Our review also emphasizes the opportunity of extracting implications for QCD axion models from primordial GW backgrounds at 3G detectors.
@@ II.1 SGWB detection methods
namely the sum of the squared SNR of each detector pair Hazboun et al. (2019a). Note that the assumption of uncorrelated noise in $N$ detectors applies to PTA experiments [^4] This assumption applies well to intrinsic pulsar noises. However, some noise sources, such as clock, planetary ephemeris errors or poorly modeled solar wind, could imprint common correlated noises in pulsar data Tiburzi et al. (2015)., where each pulsar can be considered as a detector. It does not apply to the three co-located interferometers composing the Einstein Telescope, which we treat separately below. Note also that the SNR scales with the square root of the observing time $t_{\mathrm{obs}}$.
@@ II.2.1 Earth-based detectors
Regarding ET, we work with the current baseline configuration consisting of a triangular arrangement of three Michelson interferometers with 10km arms, situated underground to minimize seismic noise. Since the three V-shape interferometers are co-located, one cannot actually use Eq. (11), as the noises will be partially correlated among the three interferometers. To take into account these partial correlations, we follow a similar approach to Branchesi et al. (2023) to compute the effective strain noise PSD. Note, however, that our resulting sensitivity differs slightly from the one calculated in Branchesi et al. (2023), see Appendix B for more details. In brief, we take the noise power spectral density $S_{\mathrm{ET-L}}$ for one ET L-shape interferometer from ET- , including a factor of $\sin^{2}(\pi/3)$ due to the triangular shape. We then use an excess cross-correlation power statistic between the three interferometers pairs, assuming a 20% correlation amongst interferometers as in Branchesi et al. (2023), which assumes a perfect characterization of the correlated noise. This constitutes a major experimental challenge that will have to be addressed to achieve the best possible sensitivity. In the weak signal regime (see Appendix B), keeping only the linear terms in the noise correlation coefficient $\Gamma^{(n)}_{12}=0.2$ (see Eq. 131), we then find
@@ II.2.2 Space-based detector
The data employed for GW search in LISA will be obtained from Time Delay Interferometry (TDI). This mitigates laser frequency noise by delaying and recombining individual measurements along spacecraft links (for further details, see Prince et al. (2002); Tinto and Dhurandhar (2021)). This process yields three Michelson TDI channels, often denoted as $X$, $Y$, and $Z$, obtained by combining measurements in each pair of arms. Assuming that the length of the arms is constant in time, it is then possible to construct uncorrelated combinations of these variables which are the $A$, $E$ and $T$ channels Estabrook et al. (2000); Vallisneri and Galley (2012). Clearly one cannot cross-correlate the output from these three channels to search for an isotropic SGWB, as the signal as well as the noise are uncorrelated amongst the channels. Furthermore, the $T$ combination, which corresponds to a Sagnac interferometer, is insensitive to the GW signal at low frequency Hogan and Bender (2001). It has been proposed to use the auto-correlation of the $A$ and $E$ channels to look for excess power due to the presence of a SGWB. [^6] Many other methods are proposed in the literature, see Boileau et al. (2021) and references therein. As explained earlier, this method is not optimal since it assumes that we have a perfect estimation of the noise in the $A$ and $E$ channels. As a result, LISA’s ability to detect cosmological backgrounds will strongly depend on the precision of noise characterization across its sensitivity frequency band. Thanks to the uncorrelated nature of these two combinations, one obtains for the total GWB SNR in LISA
@@
The detection of cosmological GWB with LISA will be affected by astrophysical foregrounds. The first we consider originates from the galactic population of white dwarf binaries. An estimation of the resulting foreground PSD after (perfect) substraction of individual bright binaries can be found in Babak et al. (2021); Karnesis et al. (2021). The level of this noise is expected to exceed the detector noise around 1mHz. Consequently, we combine the PSD of this foreground with that of the detector in quadrature. The second foreground we consider is due to the extra galactic population of unresolvable white dwarf binaries and stellar-mass black hole (BH) binaries. We model the former using Eq. (25) of Staelens and Nelemans (2024) and the latter using Eq. (20), with $\Omega_{A}=7.87\cdot 10^{-13}$ and $f_{*}=0.003$ Hz Babak et al. (2023). In the same way as for the Earth-based detectors, we compute the SNR of this foreground using Eq. (21). We find $\mathrm{SNR}^{\mathrm{LISA}}_{\mathrm{BBH+WDB}}\simeq{448}$, which is dominated by the population of extra galactic WDB since $\mathrm{SNR}^{\mathrm{LISA}}_{\mathrm{BBH}}\simeq{56}$ in agreement with Pieroni and Barausse (2020).
@@ II.2.3 Pulsar Timing Arrays
Regarding the computation of the effective strain noise PSD of SKA, we employ Eq. (11) using the hasasia Python package to produce realistic sensitivity curves Hazboun et al. (2019b, a). Indeed, PTA experiments are not only limited by the white noise from the radio telescopes and pulse jitter Goncharov et al. (2020), but also by three main components: (i) the uncertainty in the timing model parameters used to construct the expected times of arrival of the radio pulses Van Haasteren et al. (2009); Blandford et al. (1984), (ii) the potential intrinsic red noise that may impact the stability of pulsar spin at low frequencies Shannon and Cordes (2010), and (iii) the chromatic noise induced by electron density fluctuations in the interstellar medium along the path of the radio pulses to Earth Jones et al. (2017); Lam et al. (2016). The latter can, in principle, be mitigated using multi-band observation but it is still anticipated to affect the sensitivity of the experiment Janssen et al. (2015). For (i) we consider the pulsar spin rate, its derivative, its sky position and distance in the timing model, see Hazboun et al. (2019a) for their effects on the PTA sensitivity. For (ii) and (iii), using the results from currently observed millisecond pulsars Agazie et al. (2023d); Antoniadis et al. (2023c), we add an intrinsic achromatic red noise component to one-third of the pulsars in the array. We parameterize this noise by a power law PSD with an amplitude (at the reference frequency of 1/year) and spectral index, in terms of characteristic strain, of $\log_{10}A=-13.5$ and $\alpha=1/2$. This simplification highlights our current uncertainty regarding the intrinsic noise spectra and the number of SKA pulsars that will be significantly affected by non-GW noise, and, consequently, its ability to probe cosmological backgrounds.
@@ III Primordial sources: overview on first order phase transition
During a PT, the order parameter (generally a scalar field in the cosmological case) changes from a metastable vacuum into a more energetically favoured one. If the system evolves maintaining thermal equilibrium, as predicted by the SM for both the QCD Aoki et al. (2006); Stephanov (2006) and the EW PTs Gurtler et al. (1997); Aoki (1997); Kajantie et al. (1996a, b); Laine and Rummukainen (1999), no significant GW emission is expected. While GWs are indeed produced by a plasma in equilibrium, their amplitude is too small to be of observational relevance if the plasma is the one of the SM Ghiglieri et al. (2020), and even when hidden sectors are present Drewes et al. (2023) (note, however, that the question remains open in the context of thermal inflation Klose et al. (2022a, b)). On the other hand, if the PT proceeds via tunnelling or thermal fluctuations over a potential barrier Linde (1983); Hogan (1983); Turner et al. (1992), bubbles of the new phase are nucleated and expand, and ultimately percolate to complete the transition. A SGWB is then sourced by the anisotropic stresses generated by the scalar field configuration when the bubbles collide Kosowsky et al. (1992); Kosowsky and Turner (1993); Caprini et al. (2008); Huber and Konstandin (2008); Jinno and Takimoto (2017); Cutting et al. (2018, 2021), and/or by the bulk fluid motion induced by the bubble expansion and collision Witten (1984); Hogan (1986); Kamionkowski et al. (1994); Hindmarsh et al. (2014, 2015, 2017); Cutting et al. (2020); Hindmarsh and Hijazi (2019); Kosowsky et al. (2002); Dolgov et al. (2002); Caprini and Durrer (2006); Kahniashvili et al. (2008); Caprini et al. (2009b); Brandenburg et al. (2017); Niksa et al. (2018); Roper Pol et al. (2020a). Note that the phenomenology linked to the presence of bubble walls and to their collisions is very rich: for example, particles can be produced at the bubble walls, leading to additional GW production Jinno et al. (2022) and to dark matter candidates Giudice et al. (2024); and furthermore leptogenesis can occur at bubble wall collisions Cataldi and Shakya (2024). In the following, we will focus on GW production from anisotropic stresses due to the aforementioned scalar field gradients at bubble collisions, and to bulk fluid motion. GW production from topological defects, which may form when a symmetry is spontaneously broken independently of the order of the PT, is discussed in Section IV.
@@ III.1 Dynamics of a first order phase transition
The starting point to determine the PT dynamics within a given model is the effective potential of the system. In many scenarios the barrier is not present at tree level, but is it induced by radiative and/or finite-temperature corrections. An important source of uncertainty is therefore linked to the fact that a perturbative treatment of the finite temperature potential can be significantly inaccurate. The best resummation techniques are based on matching to a three-dimensional effective field theory, but they can be heavy to implement, and must be carried out on a model-by-model basis (see Athron et al. (2024) for a review and e.g. Ekstedt et al. (2023, 2024) for recent progress). This uncertainty can severely affect SGWB predictions, via the inaccurate evaluation of the PT parameters entering the GW signal Kainulainen et al. (2019); Gould et al. (2019); Niemi et al. (2019); Croon et al. (2021); Kierkla et al. (2024); Niemi et al. (2021); Schicho et al. (2021); Gould and Tenkanen (2024); Lewicki et al. (2024); Banerjee et al. (2024). Correspondingly, it can also compromise the reconstruction of the model parameters, in case a SGWB signal from a first order PT is detected Bian et al. (2020); Hashino et al. (2017); Friedrich et al. (2022); Caprini et al. (2024a).
@@ III.2 Gravitational wave generation from a first order phase transition
For the three sourcing processes described above, the tensor anisotropic stress $\Pi_{ij}=T_{ij}^{\rm{TT}}$ i.e. the transverse traceless component of the spatial part of the source energy momentum tensor $T_{ij}$, is a random variable, since the bubble nucleation process is stochastic. $\Pi_{ij}$ also has a finite lifetime $\delta\tau_{\mathrm{fin}}=\tau_{\mathrm{fin}}-\tau_{*}$, where $\tau$ denotes conformal time, we assume radiation domination and set $\tau_{*}=\mathcal{H}_{*}^{-1}=(a_{*}H_{*})^{-1}$ as the initial time of action of the source. $\delta\tau_{\mathrm{fin}}$ can be short compared to the Hubble time (e.g. for bubble collision), but can also be of several Hubble times (e.g. for turbulence). After the decay of the source, the GWs become freely propagating plane waves, redshifting with the expansion of the universe. The SGWB power spectrum today satisfies then:
@@
$h^{2}\mathrm{\Omega}_{\rm{GW}}(\tau_{0},f)=\frac{h^{2}}{\rho_{c}}\left(\frac{a_{\mathrm{fin}}}{a_{0}}\right)^{4}\rho_{\mathrm{fin}}\,\mathrm{\Omega}_{\rm{GW}}(\tau_{\mathrm{fin}},f)=h^{2}F_{\mathrm{GW},0}\,\mathrm{\Omega}_{\rm{GW}}(\tau_{\mathrm{fin}},f)\,,~{}~{}\mathrm{with}~{}~{}h^{2}F_{\mathrm{GW},0}\simeq 1.64\cdot 10^{-5}\left(\frac{100}{g_{\mathrm{fin}}}\right)^{\frac{1}{3}}.$ (42)
In order to give a compact expression for $\mathrm{\Omega}_{\rm{GW}}(\tau_{\mathrm{fin}},f)$, we assume that the GW source operates during the radiation dominated era, and is such that the total energy density in the Universe is $\rho_{\mathrm{tot}}(\tau)\propto a^{-4}(\tau)$. This assumption will always be valid in the case of weak PTs, for which the vacuum energy is subdominant throughout the PT. It is also adapted in general to sound waves and MHD turbulence, which continue to source GWs long after the bubble collisions are over and the Universe returns to a radiation dominated expansion. It is clearly not suited to the case of strong PTs, for which vacuum energy may dominate the Universe expansion for a while. However, in this case, the SGWB is mainly sourced by bubble collisions, which is a fast process occurring around percolation time. The usual assumption for this kind of source is therefore to neglect the expansion of the Universe altogether, and to assume that GW generation occurs in flat space-time.
@@
We define for later convenience the energy fraction $K(\tau)=\rho_{s}(\tau)/\rho_{\mathrm{tot}}(\tau)$, where $\rho_{s}(\tau)$ denotes the energy density of the GW sourcing process (i.e., kinetic energy of the scalar field and/or of the fluid motion), and $\rho_{\mathrm{tot}}(\tau)$ denotes the total energy density in the universe at the PT time, which we assume to be radiation. It is reasonable to expect that the source $\rho_{s}(\tau)$ also redshifts with the Universe expansion, like radiation. However, superimposed on the radiation-like scaling, one also expects a peculiar time evolution, characteristic of the source itself: for example, a growing phase followed by an overall decay. This is captured by the residual time dependence of $K(\tau)$. One can then rewrite a normalised anisotropic stress $\tilde{\Pi}_{ij}=\Pi_{ij}/(K\rho_{\mathrm{tot}})$. Further neglecting possible changes in the relativistic degrees of freedom, and averaging over the irrelevant, rapid oscillations in time of the plane waves, the GW energy density fraction when the source has stopped operating becomes then (where $k$ denotes the comoving wave-number) Caprini et al. (2008, 2009b); Caprini and Figueroa (2018)
$\mathrm{\Omega}_{\rm{GW}}(\tau_{\mathrm{fin}},k)=\frac{3}{4\pi^{2}}k^{3}\int^{\tau_{\mathrm{fin}}}_{\tau_{*}}{\frac{\mathrm{d}\tau^{\prime}}{\tau^{\prime}}}\int^{\tau_{\mathrm{fin}}}_{\tau_{*}}{\frac{\mathrm{d}\tau^{\prime\prime}}{\tau^{\prime\prime}}\cos[k(\tau^{\prime}-\tau^{\prime\prime})]\,K^{2}(\tau^{\prime},\tau^{\prime\prime})\,P_{\tilde{\Pi}}(k,\tau^{\prime},\tau^{\prime\prime})}\,,$ (43)
@@
where $P_{\tilde{\Pi}}(k,\tau^{\prime},\tau^{\prime\prime})$ is the unequal time correlator (UETC) of the normalised anisotropic stress $\tilde{\Pi}_{ij}(\tau)$ in Fourier space,
$\langle\tilde{\Pi}_{ij}(\mathbf{k},\tau^{\prime})\,\tilde{\Pi}_{ij}^{*}(\mathbf{q},\tau^{\prime\prime})\rangle=(2\pi)^{3}\,\delta^{(3)}(\mathbf{k}-\mathbf{q})\,P_{\tilde{\Pi}}(k,\tau^{\prime},\tau^{\prime\prime})\,.$ (44)
In Eq. 43, the radiation-like dependence of the GW source has been absorbed in the factors $1/\tau^{\prime}$ and $1/\tau^{\prime\prime}$, while $K^{2}(\tau^{\prime},\tau^{\prime\prime})$ contains a possible overall time dependence of the source (such as an overall growth and/or decay), and $P_{\tilde{\Pi}}(k,\tau^{\prime},\tau^{\prime\prime})$ contains the residual time dependence, typically a time decorrelation. Note that, for a source effectively operating in flat space-time such as e.g. bubble collision, one can just substitute the factors $1/\tau^{\prime}$ and $1/\tau^{\prime\prime}$ in Eq. 43 with a factor $H_{*}^{2}$ Caprini et al. (2009b).
@@
In order to find the SGWB signal from a first order PT, one therefore needs to determine the time dependence of $K^{2}(\tau^{\prime},\tau^{\prime\prime})$ together with the anisotropic stress UETC $P_{\tilde{\Pi}}(k,\tau^{\prime},\tau^{\prime\prime})$, for each of the the sourcing processes described above: bubbles and spherical fluid configurations collision, sound waves, MHD turbulence. This is a complicated task, which can be tackled both analytically and numerically. However, it turns out that, by making use of two simple models for the UETC, it is possible to understand the scaling with the sources’ parameters provided by more refined analytical and numerical evaluations. We present these models below, as an update of what discussed in sections 3.1 and 3.2 of Caprini et al. (2020). In the following we will then clarify the connection between the general scalings provided by these simple models, and what found in specific analytical and numerical evaluations on a source-by-source basis.
@@
  • The first model consists in assuming that the UETC is simply constant in time, $K^{2}(\tau^{\prime},\tau^{\prime\prime})P_{\tilde{\Pi}}(k,\tau^{\prime},\tau^{\prime\prime})=K^{2}\,P_{\tilde{\Pi}}(k)$ Roper Pol et al. (2022a). As we will see, the constant-in-time approximation can be applied to bubble collision and to (M)HD turbulence.
@@
  • The second model consists in assuming that the UETC is stationary in time, $K^{2}(\tau^{\prime},\tau^{\prime\prime})P_{\tilde{\Pi}}(k,\tau^{\prime},\tau^{\prime\prime})=K^{2}f(k,|\tau^{\prime}-\tau^{\prime\prime}|)$ Hindmarsh et al. (2014, 2015). This assumption has been used mainly for SGWB generation by sound waves (see Caprini et al. (2024b) for an alternative proposal).
    Since the source is necessarily dissipated at some point, it is “stationary” but only within its lifetime $\delta\tau_{\mathrm{fin}}$. Upon changing variable $\tau_{-}=\tau^{\prime\prime}-\tau^{\prime}$, Eq. 43 takes the form:
    $\mathrm{\Omega}_{\rm{GW}}(\tau_{\mathrm{fin}},k)\simeq\frac{3}{4\pi^{2}}K^{2}k^{3}{\left[\int_{\tau_{*}}^{\tau_{\mathrm{fin}}}\frac{\mathrm{d}\tau^{\prime}}{\tau^{\prime}}\int_{\tau_{*}-\tau^{\prime}}^{\tau_{\mathrm{fin}}-\tau^{\prime}}\mathrm{d}\tau_{-}\frac{\cos(k\tau_{-})f(k,\tau_{-})}{\tau^{\prime}+\tau_{-}}\right]\,.}$ (49)
@@
    $\mathrm{\Omega}_{\rm{GW}}(\tau_{\mathrm{fin}},k)\simeq\frac{3}{4\pi^{2}}K^{2}k^{3}\int_{\tau_{*}}^{\tau_{\mathrm{fin}}}{\frac{\mathrm{d}\tau^{\prime}}{(\tau^{\prime})^{2}}}\int_{-\infty}^{+\infty}\mathrm{d}\tau_{-}\cos(k\tau_{-})f(k,\tau_{-})\simeq\frac{3}{4\pi^{2}}K^{2}\frac{\mathcal{H}_{*}\delta\tau_{\mathrm{fin}}}{1+\mathcal{H}_{*}\delta\tau_{\mathrm{fin}}}[\mathcal{H}_{*}k^{3}F(k)]\,,$ (50)